# On Conjecture of Binomial Edge Ideals of Linear Type

### v2

@@ 5. Counterexample
Verification of non-$p$-sequence property in Remark 2.1(2).
@@
11-- Display the Quadratic relation element
12
13I_35
14
15-- Verify that the evaluation on the element a_35 yields 0
16
17(((x_4*x_8*y_3*y_7)-(x_3*x_8*y_4*y_7)-(x_4*x_7*y_3*y_8)+(x_3*x_7*y_4*y_8))*(x_6*y_2-x_2*y_6)*(-x_1*y_5+x_5*y_1))+(((x_6*x_8*y_4*y_7)-(x_4*x_8*y_6*y_7)-(x_6*x_7*y_4*y_8)+(x_4*x_7*y_6*y_8))*(-x_2*y_3+x_3*y_2)*(-x_1*y_5+x_5*y_1))+(((x_3*x_8*y_5*y_7)-(x_5*x_8*y_3*y_7)-(x_3*x_7*y_5*y_8)+(x_5*x_7*y_3*y_8))*(x_6*y_2-x_2*y_6)*(-x_1*y_4+x_4*y_1))+(((x_5*x_8*y_6*y_7)-(x_6*x_8*y_5*y_7)-(x_5*x_7*y_6*y_8)+(x_6*x_7*y_5*y_8))*(-x_2*y_3+x_3*y_2)*(-x_1*y_4+x_4*y_1))+(((x_3*x_7*y_5*y_6)-(x_3*x_6*y_5*y_7)-(x_3*x_5*y_6*y_7)+(x_5*x_6*y_3*y_7))*(-x_1*y_2+x_2*y_1)*(x_8*y_4-x_4*y_8))+(((x_4*x_6*y_5*y_8)-(x_4*x_8*y_5*y_6)-(x_5*x_6*y_4*y_8)+(x_4*x_5*y_6*y_8))*(x_7*y_3-x_3*y_7)*(-x_1*y_2+x_2*y_1))+(((x_7*x_8*y_5*y_6)-(x_6*x_8*y_5*y_7)-(x_5*x_7*y_6*y_8)+(x_5*x_6*y_7*y_8))*(-x_3*y_4+x_4*y_3)*(-x_1*y_2+x_2*y_1))
Acknowledgement. The first author thanks A. V. Jayanthan for his valuable comments and for pointing out an error in an earlier version. The second author gratefully acknowledges partial funding received from the Science and Engineering Research Board (SERB), India, under the Core Research Grant (CRG/2023/007668).

### v3

@@ 5. Counterexample
Verification of non-$p$-sequence property in Example 3.
@@
11
12-- Verify that the evaluation on the element I_35 yields 0
13
14(((x_4*x_8*y_3*y_7)-(x_3*x_8*y_4*y_7)-(x_4*x_7*y_3*y_8)+(x_3*x_7*y_4*y_8))*(x_6*y_2-x_2*y_6)*(-x_1*y_5+x_5*y_1))+(((x_6*x_8*y_4*y_7)-(x_4*x_8*y_6*y_7)-(x_6*x_7*y_4*y_8)+(x_4*x_7*y_6*y_8))*(-x_2*y_3+x_3*y_2)*(-x_1*y_5+x_5*y_1))+(((x_3*x_8*y_5*y_7)-(x_5*x_8*y_3*y_7)-(x_3*x_7*y_5*y_8)+(x_5*x_7*y_3*y_8))*(x_6*y_2-x_2*y_6)*(-x_1*y_4+x_4*y_1))+(((x_5*x_8*y_6*y_7)-(x_6*x_8*y_5*y_7)-(x_5*x_7*y_6*y_8)+(x_6*x_7*y_5*y_8))*(-x_2*y_3+x_3*y_2)*(-x_1*y_4+x_4*y_1))+(((x_3*x_7*y_5*y_6)-(x_3*x_6*y_5*y_7)-(x_3*x_5*y_6*y_7)+(x_5*x_6*y_3*y_7))*(-x_1*y_2+x_2*y_1)*(x_8*y_4-x_4*y_8))+(((x_4*x_6*y_5*y_8)-(x_4*x_8*y_5*y_6)-(x_5*x_6*y_4*y_8)+(x_4*x_5*y_6*y_8))*(x_7*y_3-x_3*y_7)*(-x_1*y_2+x_2*y_1))+(((x_7*x_8*y_5*y_6)-(x_6*x_8*y_5*y_7)-(x_5*x_7*y_6*y_8)+(x_5*x_6*y_7*y_8))*(-x_3*y_4+x_4*y_3)*(-x_1*y_2+x_2*y_1))
Acknowledgement. The first author thanks A. V. Jayanthan for his valuable comments. The first author thank the Scientific and Technological Research Council of Turkey - TÜBİTAK (Grant No: 124F113) for the financial support. The second author gratefully acknowledges partial funding received from the Science and Engineering Research Board (SERB), ANFR, India, under the Core Research Grant (CRG/2023/007668).
Competing Interests. The authors declare that they have no known competing financial interests or personal relationships that could have appeared to influence the work reported in this paper.